# Sociodynamics-inspired Adaptive Coalition and Client Selection in Federated Learning

### v1

@@ 3 FedCVR-Bolt Algorithm
First, let us formalize the underlying network structure aimed to study. Let $\theta(t)$ indicate the vector of current models at round $t$, such that $\theta_{k}(t)$ is the model associated to client $k$ at communication round $t$. As introduced in Mas2010_opinion_exp a natural way to implement homophily it to define social influence of agent $j$ on agent $k$ as
@@
where $\gamma>0$ is a parameter which amplifies the role of models similarities, i.e., the larger $\gamma$, the more pronounced is the homophily effect.
@@
Move now the focus on the coalitions formed which may vary with respect to the previous round. Formally, the goal of this step is to identify the $P$ coalitions formed at round $t$ that will be indicated by $\mathcal{C}_{1},\dots,\mathcal{C}_{P}\subset\Pi(K)$, where $\Pi(K)$ denotes the set of all possible partitions of $[K]$. Let us specify that the number of communities $P$ is not an output of the clustering step but it is fixed a priori and it coincides with the model participation rate. As proposed by Narantsatsralt2017_spectralClust_OD , a natural way to detect communities in social networks is spectral clustering. The idea is now to project the data, collected in matrix $W$, in the corresponding eigenvector space and then to properly apply a k-means on this space. The projection on eigenvectors space is crucial since it allows to significantly distinguish the similar nodes into more distanced positions in feature space, assigning similar values to members of the same community and capturing also soft boundaries between communities, both convex and non-convex.
@@
Once the $P$ coalitions of clients have been identified, for each coalition our goal is now to identify the optimal client to sample at round $t$.
Let us introduce the value vectors $v^{d}\in\mathbb{R}^{K}$ for the $d$-th model component, where each entry is defined as
$v_{k}^{d}=\dfrac{(C^{d}\alpha)_{k}^{2}}{C^{d}_{kk}}\quad\text{for all}\;k\in K\,.$ (5)
As it will be extensively explained in Section 3.1, $v_{k}^{d}$ refers to the variance reduction for the $d$-th component of global model $\theta_{gl}^{d}$ estimation subjected to $\theta_{k}^{d}$ sampling raineri2025_ECC_ext , i.e.,
$v_{k}^{d}=\mathbb{V}ar(\theta_{gl}^{d})-\mathbb{E}[\mathbb{V}ar(\theta_{gl}^{d}|\theta_{k}^{d})]\,.$
@@
$v_{k}:=\sum_{d=1}^{D}v^{d}_{k},$ (6)
which is a natural choice johnson2002_ult_statistical_anal for the overall residual variance for the global model $\theta_{gl}$, see Section 3.1 for theoretical details.
Within the defined setting, we can now properly define the sampling strategy to adopt, i.e. we select for each coalition the client maximizing vector $v$, i.e., for each coalition $p=1,\dots,P$ we select the client $j_{p}$ such that
$j_{p}\in\arg\max_{k\in\mathcal{C}_{p}(t)}v_{k}\,.$ (7)
In real case scenario, the optimization in (7) becomes overly restrictive due to the unavailability of the true data distribution, which introduces an inherent estimation error in the covariance matrix. To mitigate the impact of this uncertainty while still exploiting the informative content of the estimated variance, we incorporate an additional exploration term into the objective.
@@
We adopt a probabilistic sampling policy inspired by the principles of statistical mechanics, specifically the Boltzmann distribution gallavotti2013statistical . This approach introduces a controlled stochasticity that helps mitigate the risk of becoming trapped by early biased or noisy estimates of the covariance matrix.
Drawing an analogy from statistical mechanics gallavotti2013statistical , wherein the probability of a system occupying a specific energy state is proportional to the exponential function of the negative energy divided by temperature, we conceptualize the probability of selecting a client $k$ within coalition $\mathcal{C}_{p}(t)$ during round $t$ based on its associated value $v_{k}$ as an indicator of the desirability of sampling client $k$. Higher values signify an increased potential for variance reduction in the estimation of the global model.
The Boltzmann-like probability measure is formally defined as:
@@
where $\pi_{p}(k;t)$ is the probability that client $k$ in coalition $\mathcal{C}_{p}(t)$ is selected at round $t$. The parameter $\beta$, analogous to the inverse temperature in statistical mechanics, controls the level of exploration. In our experiments, we set $\beta=1$. This formulation guarantees that clients with higher estimated variance reduction $v_{k}$ are accorded proportionally greater probabilities of selection, thereby reflecting the scale of their potential contributions within their respective coalitions. Moreover, this framework enables the exploration of clients that, while seemingly less promising, may nonetheless yield significant benefits in contexts characterized by uncertainty.
@@
with $\rho_{kj_{p}}$ the Pearson correlation coefficient between models of client $k$ and $j_{p}$. As widely explained in Section 3.1, this is a natural choice since it coincides with taking the expected value of $\theta_{k}$ conditioned to observation $\theta_{j_{p}}$, i.e., $\mathbb{E}[\theta_{k}|\theta_{j_{p}}(t+1)]$. Precisely, the rescaling parameter coincides with the optimal regression coefficient of $\theta_{k}$ given the sample $\theta_{j_{p}}(t+1)$ under the non restrictive assumption of normalized models.
@@
Using the $k,j$-th entry of the covariance matrix, defined in (10), it is possible to obtain an online estimate of the covariance matrix, using the Robbins-Monro estimation robbins1951stochastic , as
@@
with $\bar{\theta}^{d}(t+1)$ the mean vector updated according to Equation (9).
@@ 3.1 Theoretical Derivation of FedCVR-Bolt
This section presents the core theoretical results that underpin and justify each step of the proposed algorithm. These results provide the formal grounding for our methodology. For complete proofs and additional theoretical insights, we refer the reader to Appendix A. Let us introduce the following notation, which will be utilized subsequently. Given a matrix $X\in\mathbb{R}^{K\times K}$, and given $\mathcal{A},\mathcal{B}\subseteq[K]$, we indicate with $X_{\mathcal{AB}}$ the submatrix of $X$ having rows in $\mathcal{A}$ and columns in $\mathcal{B}$.
@@
Finding the optimal subset of clients to sample in order to obtain the best estimate of the global model is an instance of the more general and widely studied subset selection problem which is known to be NP-complete kempe2008_subsetselection . Generalizing raineri2025_ECC_ext , let us here reformulate the main results on the specific context of interest. The first step consists in defining the proper evaluation metric. Specifically, we focus on Variance Reduction which measures the reduction in uncertainty on the variable to estimate, i.e., $\theta_{gl}$, conditioned to the observation done. Formally, given an arbitrary subset of clients $\mathcal{A}\subseteq\mathcal{K}$, for each model component $d$, we denote the variance reduction on $\theta_{gl}^{d}$ conditioned to the observations $\theta_{\mathcal{A}}^{d}$ as
@@
First, let us introduce a preliminary Lemma, which comes from Lemma 2 in raineri2025_ECC_ext , which finds an explicit formulation for the global model expectation conditioned to samples $\mathcal{A}\subseteq\mathcal{K}$.
@@
Consider $\theta^{d}\in\mathbb{R}^{K}$ the vector containing all the $d$-th model components of the federation and $\theta^{d}_{gl}=\alpha^{\top}\theta^{d}$ the global model to estimate. Given $\mathcal{A}\subseteq K$, it holds
This lemma is crucial, as shown in Appendix A, to prove the following core result, which generalized Proposition 2 in raineri2025_ECC_ext , which provides an explicit formulation for the computation of variance reduction function, defined in (12), applied to our specific application.
@@
Consider $\theta^{d}\in\mathbb{R}^{K}$ the vector containing all the $d$-th model components of the federation and $\theta^{d}_{gl}=\alpha^{\top}\theta^{d}$ the global model to estimate. Then, the variance reduction, subjected to sampling the subset $\mathcal{A}$ of possible clients, is computed as
@@
Consider $\theta^{d}\in\mathbb{R}^{K}$ the vector containing all the $d$-th model components of the federation and $\theta^{d}_{gl}=\alpha^{\top}\theta^{d}$ the global model to estimate. Let $\theta^{d}_{k}$ be the sampled model at round $t$. Then, the variance reduction for the global model $\theta^{d}_{gl}$ given sample $\theta_{k}^{d}$ is
$v_{k}^{d}=\frac{(C^{d}\alpha)^{2}}{C^{d}_{kk}}\,.$ (15)
Until now, we succeed in finding an explicit formulation for the variance reduction function of the global model $\theta_{gl}^{d}$ for each $d$-th component. Let us now define a collective measure of the overall variance reduction of $\theta_{gl}$. Based on johnson2002_ult_statistical_anal , we define the total variance reduction as the sum of variance reductions associated to the different model components, i.e., $v_{k}=\sum_{d=1}^{D}v_{k}^{d}$, which is the trace of $\mathbb{V}ar(\mathbb{E}(\theta_{gl}|\theta_{k}))$, which geometrically coincides with the sum of the $D$ squared deviation vectors.
@@
Once the optimal model to sample is selected through the introduced sampling strategy, we use this additional information to update all the other models distribution. Preliminary, let us recall the definition of the Pearson Correlation Coefficient, i.e., given two random variable $X_{i}$ and $X_{j}$ out of a random vector $X$ with covariance matrix $C$, it holds
$\rho_{jk}=\frac{C_{kj}}{\sqrt{C_{kk}}\sqrt{C_{jj}}}\,.$ (16)
We can now formally define the update policy for the $j$-th model given the latest $k$-th sample. For the complete proofs refer to Appendix A.
@@
Consider $\theta^{d}\in\mathbb{R}^{K}$ the vector containing all the $d$-th model components of the federation and $\theta^{d}_{gl}=\alpha^{\top}\theta^{d}$ the global model to estimate. Let $\theta^{d}_{k}$ be the sampled model at round $t$. Then, the optimal estimator of an arbitrary model $d$-th component $\theta^{d}_{j}$ with $j\neq k$ is
$\mathbb{E}[\theta_{j}^{d}|\theta_{k}^{d}]=\rho_{kj}\frac{\sqrt{v_{j}}}{\sqrt{v_{k}}}\theta_{k}^{d}\,.$ (17)
@@
Consider $\tilde{\theta}^{d}\in\mathbb{R}^{K}$ the normalized vector containing all the $d$-th model components of the federation, such that $\mathbb{V}ar(\tilde{\theta}^{d})=\mathbf{1}$. Let $\tilde{\theta}^{d}_{k}$ be the sampled model at round $t$. Then, the optimal estimator of an arbitrary model $d$-th component $\tilde{\theta}^{d}_{j}$ with $j\neq k$ is
$\mathbb{E}[\tilde{\theta}_{j}^{d}|\tilde{\theta}_{k}^{d}]=\rho_{kj}\tilde{\theta}_{k}^{d}\,.$ (18)

### v2

@@ 3 FedCVR-Bolt Algorithm Design
First, let us formalize the underlying network structure aimed to study. Let $\theta(t)$ indicate the vector of current models at round $t$, such that $\theta_{k}(t)$ is the model associated to client $k$ at communication round $t$. As introduced in (Mäs et al., 2010) a natural way to implement homophily it to define social influence of agent $j$ on agent $k$ as
@@
where $\gamma>0$ is a parameter which amplifies the role of models similarities, i.e., the larger $\gamma$, the more pronounced is the homophily effect. The resulting matrix admits a clear social-systems interpretation: agents (clients) with similar “opinions” (model parameters) exert stronger mutual influence and naturally form clusters. Importantly, computing this matrix, which is standard practice in clustered and personalized FL (Sattler et al., 2020; Bao et al., 2023), here reduces to evaluating L2 distances between server-available parameter vectors. To further validate our approach, we additionally implemented our algorithm with other similarity functions used in coalition formation (Scholkopf & Smola, 2001); see Appendix B.1.
@@
Move now the focus on the coalitions formed which may vary with respect to the previous round. Formally, the goal of this step is to identify the $P$ coalitions formed at round $t$ that will be indicated by $\mathcal{C}_{1},\dots,\mathcal{C}_{P}\subset\Pi(K)$, where $\Pi(K)$ denotes the set of all possible partitions of $[K]$. Notice that the number of communities $P$ neither is assumed to be the true number of clusters in the data, nor is an output of the clustering step, but it is fixed a priori and it coincides with the model participation rate. This assumption is not restrictive since it is a common practice in clustered framework (e.g., refer to IFCA (Ghosh et al., 2020) and FeSEM (Long et al., 2023)).
As proposed in Narantsatsralt & Kang (2017), a natural way to detect communities in social networks is spectral clustering. The idea is now to project the data, collected in matrix $W$, in the corresponding eigenvector space and then to properly apply a k-means on this space. The projection on eigenvectors space is crucial since it allows to significantly distinguish the similar nodes into more distanced positions in feature space, assigning similar values to members of the same community and capturing also soft boundaries between communities, both convex and non-convex.
@@
Once the $P$ coalitions of clients have been identified, for each coalition our goal is now to identify the optimal client to sample at round $t$. Let us introduce the value vectors $v^{d}\in\mathbb{R}^{K}$ for the $d$-th model component, where each entry is defined as
$v_{k}^{d}=\dfrac{(C^{d}\alpha)_{k}^{2}}{C^{d}_{kk}}\quad\text{for all}\;k\in\mathcal{K}\,.$ (5)
As it will be explained in Section 3.1, $v_{k}^{d}$ refers to the variance reduction for the $d$-th component of global model $\theta_{gl}^{d}$ estimation subjected to $\theta_{k}^{d}$ sampling (Raineri et al., 2025).
@@
$v_{k}:=\sum\nolimits_{d=1}^{D}v^{d}_{k},$ (6)
which is a natural choice (Johnson & Wichern, 2002) for the overall residual variance for the global model $\theta_{gl}$, see Section 3.1 for theoretical details.
Within the defined setting, we select for each coalition the client maximizing vector $v$, i.e., for each coalition $p=1,\dots,P$ we select the client $j_{p}$ such that
$j_{p}\in\arg\max\nolimits_{k\in\mathcal{C}_{p}(t)}v_{k}\,.$ (7)
Selecting one representative per cluster mitigates overfitting to any single sub-distribution while preserving global diversity. Optimizing for variance reduction captures cross-client coupling, yielding more stable updates.
In real case scenario, the optimization in equation 7 becomes overly restrictive due to the unavailability of the true data distribution, which introduces an inherent estimation error in the covariance matrix. To mitigate the impact of this uncertainty while still exploiting the informative content of the estimated variance, we incorporate an additional exploration term into the objective.
@@
Instead of a greedy, deterministic selection based on the estimated variance, we employ a Boltzmann exploration policy (Gallavotti, 2013). This introduces a controlled stochasticity that helps mitigate the risk of becoming trapped by early biased or noisy estimates of the covariance matrix (Powell, 2022). Drawing an analogy with statistical mechanics (Gallavotti, 2013), where the probability of a system occupying a specific energy state is proportional to the exponential function of the negative energy divided by temperature, we let the probability of selecting a client $k$ within coalition $\mathcal{C}_{p}(t)$ during round $t$ based on its associated value $v_{k}$ be an indicator of the desirability of sampling client $k$. Higher values indicate greater potential to reduce the variance of the global model estimate. This Boltzmann-like probability measure is formally defined as:
@@
where $\pi_{p}(k;t)$ is the probability that client $k$ in coalition $\mathcal{C}_{p}(t)$ is selected at round $t$. The parameter $\beta$, analogous to the inverse temperature in statistical mechanics, controls the level of exploration. In our experiments, we set $\beta=1$. This design assigns clients with larger estimated variance-reduction $v_{k}$ proportionally higher selection probabilities, aligning sampling with their expected contribution within each coalition. Moreover, this framework keeps nonzero probability for lower-scoring clients, preventing myopic exploitation and enabling gains under uncertainty.
@@
with $\rho_{kj_{p}}$ the Pearson correlation coefficient between models of client $k$ and $j_{p}$. As explained in Section 3.1, this is a natural choice as it coincides with taking the expected value of $\theta_{k}$ conditioned to observation $\theta_{j_{p}}$, i.e., $\mathbb{E}[\theta_{k}|\theta_{j_{p}}(t+1)]$. Precisely, the rescaling parameter coincides with the optimal regression coefficient of $\theta_{k}$ given $\theta_{j_{p}}(t+1)$ under the assumption of normalized models.
@@
Using the $k,j$-th entry of the covariance matrix in equation 10, it is possible to obtain an online estimate of the covariance matrix, using the Robbins-Monro estimation (Robbins & Monro, 1951), as
@@
with $\bar{\theta}^{d}(t+1)$ the mean vector updated according to equation 9.
@@ 3.1 Theoretical Derivation of FedCVR-Bolt
This section presents the core theoretical results that underpin and justify each step of the proposed algorithm, providing the formal grounding for our methodology. For complete proofs and additional theoretical insights, we refer the reader to Appendix B. Let us introduce the following notation, which will be utilized subsequently. Given a matrix $X\in\mathbb{R}^{K\times K}$, and given $\mathcal{A},\mathcal{B}\subseteq[K]$, we indicate with $X_{\mathcal{AB}}$ the submatrix of $X$ having rows in $\mathcal{A}$ and columns in $\mathcal{B}$.
@@
Finding the optimal subset of clients to sample in order to obtain the best estimate of the global model is an instance of the more general and widely studied subset selection problem which is known to be NP-complete (Das & Kempe, 2008). Generalizing Raineri et al. (2025), let us here reformulate the main results on the specific context of interest. The first step consists in defining the proper evaluation metric. Specifically, we focus on Variance Reduction which measures the reduction in uncertainty on the variable to estimate, i.e., $\theta_{gl}$, conditioned to the observation done. Formally, given an arbitrary subset of clients $\mathcal{A}\subseteq\mathcal{K}$, for each model component $d$, we denote the variance reduction on $\theta_{gl}^{d}$ conditioned to the observations $\theta_{\mathcal{A}}^{d}$ as
@@
First, let us introduce a preliminary Lemma, which comes from Lemma 2 in Raineri et al. (2025), which finds an explicit formulation for the global model expectation conditioned to samples $\mathcal{A}\subseteq\mathcal{K}$.
@@
Consider $\theta^{d}\in\mathbb{R}^{K}$ the vector containing all the $d$-th model components of the federation and $\theta^{d}_{gl}=\alpha^{\top}\theta^{d}$ the global model to estimate. Given $\mathcal{A}\subseteq K$, it holds [^1] Here, with some abuse of notation, $\mathbb{E}[\theta_{gl}^{d}|\theta_{\mathcal{A}}^{d}]$ denotes the projection of $\theta_{gl}^{d}$ onto the space of all linear combinations of the components of $\theta_{\mathcal{A}}^{d}$ with constant coefficients, found via linear regression. Under standard assumption of normal distribution it would coincide with the true conditional expectation (Hastie et al., 2009). $\mathbb{E}[\theta^{d}_{gl}|\theta^{d}_{\mathcal{A}}]=(C^{d}_{\mathcal{A}\mathcal{A}})^{-1}(C^{d}\alpha)_{\mathcal{A}}\theta^{d}_{\mathcal{A}}\,$, where $C^{d}_{\mathcal{A}\mathcal{A}}$ is the invertible covariance matrix of $\theta_{\mathcal{A}}^{d}.$
This lemma is crucial, as shown in Appendix B, provide an explicit formulation for variance reduction function in equation 12, properly generalizing Proposition 2 in Raineri et al. (2025).
@@
Consider $\theta^{d}\in\mathbb{R}^{K}$ the vector containing all the $d$-th model components of the federation and $\theta^{d}_{gl}=\alpha^{\top}\theta^{d}$ the global model to estimate. Then, the variance reduction, subjected to sampling the subset $\mathcal{A}$ of possible clients, is computed as $v^{d}_{\mathcal{A}}=(C^{d}\alpha)_{\mathcal{A}}^{\top}(C^{d}_{\mathcal{A}\mathcal{A}})^{-1}(C^{d}\alpha)_{\mathcal{A}}$, where $C^{d}$ is the covariance matrix which captured the correlations among the clients.
@@
Consider $\theta^{d}\in\mathbb{R}^{K}$ the vector containing all the $d$-th model components of the federation and $\theta^{d}_{gl}=\alpha^{\top}\theta^{d}$ the global model to estimate. Let $\theta^{d}_{k}$ be the sampled model at round $t$. Then, the variance reduction for the global model $\theta^{d}_{gl}$ given sample $\theta_{k}^{d}$ is $v_{k}^{d}=\frac{(C^{d}\alpha)^{2}}{C^{d}_{kk}}$.
So far, we succeed in finding an explicit formulation for the variance reduction function of the global model $\theta_{gl}^{d}$ for each $d$-th component. Let us now define a collective measure of the overall variance reduction of $\theta_{gl}$. Based on Johnson & Wichern (2002), we define the total variance reduction as the sum of variance reductions associated to the different model components, i.e., $v_{k}=\sum_{d=1}^{D}v_{k}^{d}$, which is the trace of $\mathbb{V}ar(\mathbb{E}(\theta_{gl}|\theta_{k}))$, coinciding with the sum of the $D$ squared deviation vectors.
@@
Once the optimal model to sample is selected through the sampling strategy, we use this additional information to update all the other models distribution. Preliminary, let us recall the definition of the Pearson Correlation Coefficient, i.e., given two random variables $X_{i}$ and $X_{j}$ out of a random vector $X$ with covariance matrix $C$, it holds $\rho_{jk}=\frac{C_{kj}}{\sqrt{C_{kk}}\sqrt{C_{jj}}}$. We can now define the update policy for the $j$-th model given the latest $k$-th sample. For more details see Appendix B.
@@
Let $\theta^{d}\in\mathbb{R}^{K}$ be the vector containing all the $d$-th model components of the federation and $\theta^{d}_{gl}=\alpha^{\top}\theta^{d}$ the global model to estimate. Let $\theta^{d}_{k}$ be the sampled model at round $t$. Then, the optimal estimator of an arbitrary model $d$-th component $\theta^{d}_{j}$ with $j\neq k$ is $\mathbb{E}[\theta_{j}^{d}|\theta_{k}^{d}]=\rho_{kj}\frac{\sqrt{v_{j}}}{\sqrt{v_{k}}}\theta_{k}^{d}$.
@@
Consider $\tilde{\theta}^{d}\in\mathbb{R}^{K}$ the normalized vector containing all the $d$-th model components of the federation, such that $\mathbb{V}ar(\tilde{\theta}^{d})=\mathbf{1}$. Let $\tilde{\theta}^{d}_{k}$ be the sampled model at round $t$. Then, the optimal estimator of an arbitrary model $d$-th component $\tilde{\theta}^{d}_{j}$ with $j\neq k$ is $\mathbb{E}[\tilde{\theta}_{j}^{d}|\tilde{\theta}_{k}^{d}]=\rho_{kj}\tilde{\theta}_{k}^{d}$.
4 Algorithm Performance Analysis
Algorithm Convergence Analysis
At each step every selected client $k$ performs local training iterations of a stochastic optimizer (e.g., Stochastic Gradient Descent) which ensures non-increasing local loss $\mathcal{L}_{k}$. Then, since the global loss is a linear combination of the local ones, i.e., $\mathcal{L}(\theta):=\sum_{k=1}^{K}\alpha_{k}\,\mathcal{L}_{k}(\theta)$, it is non-increasing. Our method can be seen as an instance of partial sampling strategy which is known to preserve the optimality guarantees. Details are provided in Appendix B.4
Computational Cost Computation
The main overhead lies in performing spectral clustering on the server at each communication round. The computational cost at each round can be broken down as follows (von Luxburg, 2007). First, for the similarity matrix construction, pairwise similarities between the $K$ clients are computed using a RBF kernel. This requires evaluating squared Euclidean distances between $K$ vectors in $\mathbb{R}^{D}$, resulting in a complexity of $\mathcal{O}(K^{2}D)$. Second, constructing the normalized Laplacian from the similarity matrix incurs a cost of $\mathcal{O}(K^{2})$. Third, extracting the top $P$ eigenvectors of the $K\times K$ Laplacian requires $\mathcal{O}(K^{2}P)$ operations in standard implementations. Fourth, the final step clusters the $K$ clients in the $P$-dimensional eigenspace, with cost $\mathcal{O}(KPI)$, where $I$ is the number of K-means iterations (in our setting constant set to 300). Overall, the dominant cost is $\mathcal{O}(K^{2}D+K^{2}P)$. In practice, since $P\ll D$ (e.g., $P=10$, $D=300$), the complexity is effectively driven by the similarity matrix computation, yielding a leading cost of $\mathcal{O}(K^{2}D)$ per round. This overhead is modest and entirely server-side, since server resources are not a bottleneck.